OpenStudy (anonymous):

The number of integers satisfying the inequality is- \[\log_{2}\sqrt{x} - (\log_{1/4}x^{2})^{2} + 1 > 0\] Ans.3

OpenStudy (anonymous):

Note : \[(\log_{1/4}x^{2})^{2} = \log^{2}_{1/4}x^{2}\]

OpenStudy (anonymous):

pls anyone give me solution.

OpenStudy (anonymous):

@lgbasallote Can you help??

OpenStudy (lgbasallote):

no idea sorry

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

@nitz Now,Only you could help me on OS.

OpenStudy (anonymous):

are you sure the answer is 3 because i am getting 2 i simplified the equation by changing the base to natural log then i got a quadratic equation in log with inequality solving it i got 2.429>x>0.581

OpenStudy (anonymous):

yes it is 3.

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

@nitz Help i am coming back after 30mins.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

@satellite73 THE BOSS!

OpenStudy (anonymous):

SHUBHAM THE SOLUTION IS GOING TOO LENGHTY

OpenStudy (anonymous):

@nitz tell me starting steps

OpenStudy (anonymous):

according to me \[\log_{2}\sqrt{x}=(1/2)\log_{2} (x)\]

OpenStudy (anonymous):

\[(\log_{1/4} (x ^{2}))^{2}=(2\log_{1/4}x)^{2} \]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

explain the above step as to how we arrived at it?

OpenStudy (anonymous):

which one ?

OpenStudy (anonymous):

\[\log_{b}(x ^{n})=n \log_{b} (x)\]

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where is 2?

OpenStudy (anonymous):

not 2, where is 1 from the ques.?

OpenStudy (anonymous):

1 from question????????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

matlab kya?

OpenStudy (anonymous):

guess we want to write everything with one base right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

मतलब ques. में जो 1 हे उसका क्या हुआ ?

OpenStudy (anonymous):

@nitz i got the eq, x+2x+2>0

OpenStudy (anonymous):

final one

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i am getting a quadratic in final answer

OpenStudy (anonymous):

yes me too it is - 2x^2 +x +1>0

OpenStudy (anonymous):

that is 2x^2

OpenStudy (anonymous):

without negative

OpenStudy (anonymous):

\[t-2t ^{2}+2>0\]

OpenStudy (anonymous):

i got \(2x^2+x-1<0\)

OpenStudy (anonymous):

minus plus kab bana???

OpenStudy (anonymous):

wait checking

OpenStudy (anonymous):

assuming of course that \(x=\log_{\frac{1}{4}}(x)\)

OpenStudy (anonymous):

i am assuming t=\[\log_{2}(x) \]

OpenStudy (anonymous):

but i am algebra challenged today so it could be wrong

OpenStudy (anonymous):

hello james!

OpenStudy (jamesj):

Hi sat! ************* If \[\log_{2}\sqrt{x} - (\log_{1/4}x^{2})^{2} + 1 > 0\] then writing everything with base e, \[ \frac{(1/2) \ln x}{\ln 2} - \left(\frac{2\ln x}{-2 \ln 2}\right)^2 + 1 > 0 \] Now simplify this one or two more steps.

OpenStudy (anonymous):

i used \(\log_2(x)=-2\log_{\frac{1}{4}}(x)\)

OpenStudy (jamesj):

Makes sense.

OpenStudy (anonymous):

ya i too got same as james

OpenStudy (anonymous):

works out nicely because you get \[-t-2t^2+1>0\] \[2t^2+t-1<0\] \[(2t-1)(t+1)<0\] \[-1<t<\frac{1}{2}\]

OpenStudy (anonymous):

now i am getting \[2t^{2}+2>0\]

OpenStudy (anonymous):

where \(t=\log_{\frac{1}{4}}(x)\)

OpenStudy (anonymous):

@satellite73 what values you getting after replacing t?

OpenStudy (anonymous):

you mean after \(-1<\log_{\frac{1}{4}}(x)<\frac{1}{2}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am confuse with powers.

OpenStudy (anonymous):

ok well the base here is \(\frac{1}{4}\) so this log is a strictly decreasing function we solve for \(x\) via \[(\frac{1}{4})^{-1}>x>(\frac{1}{4})^{\frac{1}{2}}\]

OpenStudy (anonymous):

@satellite73 your ans. is correct. Now steps .

OpenStudy (anonymous):

or in other words \[\frac{\sqrt{2}}{2}<x<4\]

OpenStudy (anonymous):

1/2 < x < 4

OpenStudy (anonymous):

yeah that

OpenStudy (anonymous):

now tell me steps from starting

OpenStudy (anonymous):

so i guess integer solutions are 1, 2 and 3 halas

OpenStudy (anonymous):

you want all the gory details?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first i wrote everything in terms of \(\log_{\frac{1}{4}}(x)\)

OpenStudy (anonymous):

because that is what is being squared so lets start from the beginning

OpenStudy (anonymous):

what eq. you got after writing in terms of base 1/4

OpenStudy (anonymous):

\[\log_2(\sqrt{x})=\frac{1}{2}\log_2(x)=\frac{1}{2}\times (-2\log_{\frac{1}{4}}(x))=-\log_{\frac{1}{4}}(x)\]

OpenStudy (anonymous):

\[\log_{\frac{1}{4}}(x^2)=2\log_{\frac{1}{4}}(x)\]

OpenStudy (anonymous):

replacing \(t=\log_{\frac{1}{4}}(x)\) you get \[-t-4t^2+1>0\] damn i forgot to square the two first time. ok lets see \[4t^2+t-1<0\] now my pretty solution is all messed up

OpenStudy (anonymous):

damn damn damn sam? any ideas?

sam (.sam.):

nowonder , I'm getting -t-4t^2+1>0 just now :)

OpenStudy (anonymous):

oh hell i thought this was so nice now it is a mess

sam (.sam.):

ok, -t-4t^2+1>0 quadratic equation \[t=\frac{-(1)\pm \sqrt{1-4(-4)}}{2(-4)}\] \[t=\frac{1\pm \sqrt{17}}{8}\]

OpenStudy (anonymous):

lets change to \(\log_2(x)\)

sam (.sam.):

I tried changing to log base 2, but doesn't work out \[4\log_2 \sqrt{x}-(\log_2x^2)^2+4=0\]

OpenStudy (anonymous):

\[\frac{1}{2}\log_2(x)-(-\log_2(x))^2+1>0\]

OpenStudy (anonymous):

damn damn damn

OpenStudy (anonymous):

ok same thing either way

OpenStudy (anonymous):

well i messed this all up now i am wondering what to do to get the answer

OpenStudy (anonymous):

now i am only getting two integer solutions, 1 and 2

OpenStudy (anonymous):

log(sub 2)(sqrt(x))-(log(sub 1/4)(x^2))^2+1>0 ln(sqrt(x))/ln(2)-(ln(x^2)/ln(1/4))^2+1>0 ln((x)^1/2)/ln(2)-(2ln(x)/ln(1/4))^2+1>0 ln(x)/ln(4)-4ln^2(x)/ln(1/4)^2+1>0 ln(x)/ln(4)-4ln^2(x)/ln(1/4)^2+1>0 ln(x)=t At+Bt^2+C>0 A=1/ln(4)=0.72 B=-4/ln(1/4)^2=-2.08 C=1 t=0.88772=ln(x)=2.429 t=-0.5415=ln(x)=0.581 What about this?

sam (.sam.):

I got it, from \[t=\frac{-1\pm \sqrt{17}}{8}\] \[\log_{1/4}x=\frac{-1\pm \sqrt{17}}{8}\] \[\huge x=\frac{1}{4}^{\frac{(-1\pm \sqrt{17})}{8}}\] ------------------------

OpenStudy (anonymous):

then................

sam (.sam.):

I missed a minus sign in the exponent just now, lol

sam (.sam.):

@shubham.bagrecha that's it

OpenStudy (anonymous):

yeah that is it, but with this weird answer there are only two solutions i am going to make a guess that there was an error somewhere in the problem. maybe it should not have been \(\log_{\frac{1}{4}}(x^2)\) but rather \(\log_{\frac{1}{4}}(x)\) which would make the solution much nicer and in fact the three integer solutions would be 1, 2, 3

OpenStudy (anonymous):

finally i got the answer.